OpenStudy (anonymous):

Help me solve for z y and z!! Please show work and my teacher said something about making it into a matrix? x-3y+2z=11 -x+4y+3z=5 2x-2y-4z=2

OpenStudy (anonymous):

I wanted to solve it by substitution or elimination, thats how we learned it in class. And im in Algebra II

OpenStudy (anonymous):

thank you kind sir.

OpenStudy (rogue):

Since you have 3 equations and 3 variables, the system is also solvable by substitution. I never learned matrices though, it might be faster. ;)

OpenStudy (anonymous):

Ty, but how would you use substitution to find each? I know how to do it for 2 variables, but not 3. My teacher doesn't explain things very well in school.

OpenStudy (rogue):

It's kind of the same process. You first solve for one variable in the first equation and substitute it into the 2nd and 3rd to make them have 2 variables. Then you solve for another variable in the new 2nd equation and substitute that into the 3rd, to get a solvable equation with one variable. Next you just solve for that 3rd variable and use it to find your first two.

OpenStudy (anonymous):

I'm not really following, can you use this question ad an example and show your work?

OpenStudy (anonymous):

add first two equations you will have equation in two variables y and z . y+5z=16 multiply the second equation (-x+4y+3z=5) with 2 and add to the third equation 6y+2z=12 solve these two equations for y and z y+5z=16 6y+2z=12 can you solve these ??

OpenStudy (anonymous):

ok i got to y+5z=16. lemme keep going...

OpenStudy (anonymous):

wait why multiply by 2

OpenStudy (rogue):

x - 3y + 2z = 11 -x + 4y + 3z = 5 2x - 2y - 4z = 2 Well first you can solve for x in your first equation. x - 3y + 2z = 11 x = 11 + 3y - 2z Next you can substitute that x into your second equation. -x + 4y + 3z = 5 -(11 + 3y - 2z) + 4y + 3z = 5 y + 5z = 16 Solve for y. y + 5z = 16 y = 16 - 5z Substitute for x & y in your third equation and solve for z. 2x - 2y - 4z = 2 2(11 + 3y - 2z) - 2(16 - 5z) - 4z = 2 Then just use your z value to find x & y.

OpenStudy (rogue):

Or you can also skip some of that substitution using Sami's clever manipulations :)

OpenStudy (anonymous):

I'll be right back. I need cake. But im prepared for an all-nighter

OpenStudy (anonymous):

i would say make your life easir with following. add first two equations you will have equation in two variables y and z . y+5z=16 multiply the second equation (-x+4y+3z=5) with 2 and add to the third equation 6y+2z=12 solve these two equations for y and z y+5z=16 ............ 1 6y+2z=12 ............. 2 from equation 1 above y=16-5z put in equation 2 6(16-5z)+2z=12 96-30z+2z=12 96-28z=12 28z=84 z=84/28=3 since y=16-5z=16-5(3)=1 z=3 y=1 put in any of the following equations to get x . x-3y+2z=11 -x+4y+3z=5 2x-2y-4z=2

OpenStudy (anonymous):

Thank you so mush both of you (:

OpenStudy (anonymous):

wait now im confused again :/

OpenStudy (tkhunny):

\[\left[\begin{matrix}1 & -3 & 2 \\ -1 & 4 & 3 \\ 2 & -2 & -4\end{matrix}\right]\] It would help if you could find the inverse matrix.

OpenStudy (rogue):

Are you still working on this? x - 3y + 2z = 11 -x + 4y + 3z = 5 2x - 2y - 4z = 2 If so, what are confused about?

OpenStudy (anonymous):

Use the following to check you answers: {x = 8, y = 1, z = 3}